At one high school, the mean time for running the 100-yard dash is 15.2 seconds with a standard deviation of 0.9 seconds. The time are very closely approximated by a normal curve. Find the percent of times that are:
More than 15.2 seconds
Between 14.3 and 16.1 seconds
Thank you.

Answer by stanbon(75887)

At one high school, the mean time for running the 100-yard dash is 15.2 seconds with a standard deviation of 0.9 seconds. The time are very closely approximated by a normal curve. Find the percent of times that are:
More than 15.2 seconds
z(15.2) = (15.2-15.2)/0.9 = 0
P(z > 15.2) = 0.50 = 50%
--------------------------------
Between 14.3 and 16.1 seconds
z(14.3) = (14.3-15.2)/0.9 = -1
z(16.1) = (16.1-15.2)/0.9 = 1
--------------------------------
P(14.3 < x < 16.1) = P(-1< z < 1) = 0.6827
===============================================
cheers,
Stan H.